OpenStudy (anonymous):

What is the gcf of 48 80

OpenStudy (dangerousjesse):

Divide each one by one, two, three, four, five, etc, until you find the greatest common divisor. GCD=GCF

jhonyy9 (jhonyy9):

48=2*2*2*2*3 80=2*2*2*2*5 --------------- gcf = 2*2*2*2 = 16 ok ?

OpenStudy (anonymous):

Thanks

jhonyy9 (jhonyy9):

was my pleasure

jhonyy9 (jhonyy9):

hope you understand it bye

OpenStudy (anonymous):

What is gcf of 85 and 51